OpenStudy (jazilove):

Pleast help fast! WILL MEDAL! The sum of the angle measures of a polygon with s sides is 2,340 degrees. Find s.

OpenStudy (welshfella):

use the formula 180(n - 2) = 2340 where n is the number of sides you require

OpenStudy (welshfella):

solve for n

OpenStudy (jazilove):

That's what I'm having trouble with.

OpenStudy (welshfella):

divide both sides by 180:- n - 2 = 2340/180 now its easy to solve

OpenStudy (jazilove):

n=15 ?

OpenStudy (welshfella):

divide 2340 by 180 yes you got it

OpenStudy (jazilove):

thank you(:

OpenStudy (welshfella):

yw